Case 5 Auburn Circular Club Pro Rodeo Roundup Developed by Jess ica Jonson Frazier Eastern Kentucky University, and Patricia H. Mounce, University of Central Arkansas The Business Situation When she was asked t Shelley Jones became president-elect of the Circular Club of Auburn, Kansas o suggest a new fundraising activity for the club. After a consider- roposed that the Circular Club sponsor a pro- amount of research, Shelley p fessional rodeo. In her presentation to the club, Shelley said that she wanted a fundraiser that would (1) continue to get better each year, (2) give back to the com- munity, and (3) provide the club a presence in the community. Shelleys goal was to have an activity that would become an annual community event and that would break even the first year and raise $5,000 the following year: In addition, based on the experience of other communities, Shelley believed that a rodeo could grow in popularity so that the club would eventually earn an average of $20,000 annually. A rodeo committee was formed. Shelley contacted the worlds oldest and largest rodeo-sanctioning agency to apply to sponsor a professional rodeo. The sanction- ing agency requires a rodeo to consist of the following five events: Bareback Riding, Bronco Riding, Steer Wrestling, Bull Riding, and Calf Roping. Because there were a number of team ropers in the area and because they wanted to include females in the competition, members of the rodeo committee added Team Roping and Womens Barrels. Prize money of $3,000 would be paid to winners in each of the seven events Members of the rodeo committee contracted with RJ Cattle Company, a livestock contractor on the rodeo circuit, to provide bucking stock, fencing, and chutes. Realizing that costs associated with the rodeo were tremendous and that ticket sales would probably not be sufficient to cover the costs, the rodeo com- mittee sent letters to local businesses soliciting contributions in exchange for various sponsorships. Exhibiting Sponsors would contribute $1,000 to exhibit their products or services, while Major Sponsors would contribute $600. Chute Sponsors would contribute $500 to have the name of their business on one of the six bucking chutes. For a contribution of $100, individuals would be included in a Friends of Rodeo list found in the rodeo programs. At each performance the rodeo announcer would repeatedly mention the names of the businesses and in- dividuals at each level of sponsorship. In addition, large signs and banners with the names of the businesses of the Exhibiting Sponsors, Major Sponsors, and Chute Sponsors were to be displayed prominently in the arena.case 5 Cases for Management Decision-Making divide the profits with the Circular Club. The Auburn Circular Club Pro Rodeo Roundup would be held on June 1, 2, and 3. The cost of an adult ticket was set at $8 in advance or $10 at the gate; the cost of a ticket for a child 12 or younger was set at $6 in advance or $8 at the gate. Tickets were not date-specific. Rather, one ticket would admit an individual to one performance of his or her choice- Friday, Saturday, or Sunday. The rodeo committee was able to secure a location through the county supervisors board at a nominal cost to the Circular Club. The atrangement allowed the use of the county fair grounds and arena for a one-week period. Several months prior to the rodeo, members of the rodeo committee had en assured that bleachers at the arena would hold 2,500 patrons. On Saturday night, paid attendance was 1,663, but all seats were filled due to poor gate con- A local youth group was contacted to provide concessions to the public a trols. Attendance was 898 Friday and 769 on Sunday. The follo wing revenue and expense figures relate to the first year of the roded illustrat Revenue data, ye Contributions from sponsors $22,000 28,971 Receipts from ticket sales Share of concession profits Sale of programs 600 Total receipts $53,084 Expenses Livestock contractor 26,000 21,000 Contestant hospitality Sponsor signs for arena 1,900 1,800 1,050 Insurance Ticket printing Sanctioning fees Entertainment 859 750 Judging fees Port-a-potties Rent Hay for horses 600 538 500 450 325 300 Western hats to first 500 children Hotel rooms for stock contractor Utilities Sand for arena Miscellaneous fixed costs Tota 105 61,410 $(8.326) Net loss *The club contracted with a local caterer to provide a tent and food for the contestants. The cost of the food was contingent on the number of contestants each evening. Information con cerning the number of contestants and the costs incurred are as follows: Friday Saturday Sunday Contestants 68 96 83 Total Cost $ 998 1.100 $3,341 On Wednesday after the rodeo, members of the rodeo committee met to dis- cuss and critique the rodeo. Jonathan Edmunds, CPA and President of the Cir- cular Club, commented that the club did not lose money. Rather, Jonathan said, The club made an investment in the rodeo.case S Cases for Management Decision-Making I. Do you think it was necessary for Shelley Jones to stipulate that she wanted a fundraiser 2. What did Jonathan Edmunds mean when he said the club had made an investment in 3. Is Jonathans comment concerning the investment consistent with Shelleys idea th structions Answer each of the following questions. that would (I) continue to get better each year; (2) give back to the community, and (3) pro- vide the club a presence in the community? Why or why not? the club should have a fundraiser that would (1) continue to get better each year, (2) give back to the community, and (3) provide the club a presence in the community? the Why or why not? What do you believe is the behavior of the rodeo expenditures in relation to ticket sales? low method. 5. Determine the fixed and variable cost components of the catering costs using the high- 6. Assume you are elected chair of the rodeo committee for next year: What steps would 7. Shelley Jonathan, and Adrian Stein, the Fundraising Chairperson, are beginning to make 4. you suggest the committee take to make the rodeo profitable? plans for next years rodeo. Shelley believes that by negotiating with local feed stores, inn- carefully analyzing costs, Jonathan has estimated that the fixed expenses can be pared keepers, and other business owners, costs can be cut dramatically. Jonathan agrees. After approximately $51,000. In addition, Jonathan estimates that variable costs are 4% of Iter talking with business owners who attended the rodeo, Adrian is confident rom sponsors at $25,600. The local youth group is unwilling to provide con- to total gross receipts. that funds solicited from sponsors will increase. Adrian is comfortable in bud cessions to the audience unless they receive all of the profits. Not having the personnel to staff the concession booth, members of the Circular Club reluctantly agree to let th uth group have 100% of the profits from the concessions. In addition, members o the rodeo committee, recognizing that the net income from programs was only $100 decide not to sell rodeo ticket sales assuming Adrian and Jonathan are correct in their assumptions. programs next year. Compute the break-even point in dollars of 8. Shelley has just learned that you are calculating the break-even point in dollars of ticket sales. She is still convinced that the Club can make a profit using the assump- tions in number 7 above (a) Calculate the dollars of ticket sales needed in order to earn a target profit of (b) Calculate the dollars of ticket sales needed in order to earn a target profit of $6,000 $12,000 9. Are the facilities at the fairgrounds adequate to handle crowds needed to generate ticket revenues calculated in number 8 above to earn a $6,000 profit? Show calcula tions to support your answers. 10. Prepare a budgeted income statement for next year using the estimated revenues from sponsors and other assumptions in number 7 above. In addition, use ticket sales based on the target profit of $12,000 estimated in 8(b). The cost of the livestock contractor, prize money, sanctioning fees, entertainment, judging fees, rent, and utilities will re main the same next year. Changes in expenses include the following: Members of the Club have decided to eliminate all costs related to contestant hospitality by soliciting a tent and food for the contestants and taking care of the Contestant Hospitality Tent themselves. The county has installed permanent restrooms at the arena, eliminating the need to rent port potties. The rodeo committee intends to pursue arrangements to have hotel rooms, hay and childrens hats provided at no charge in exchange for sponsorships. The cost of ban- ners varies with the number of sponsors. Signs and More charged the Circular Club $130 for each Exhibiting Sponsor banner and $48 for each Major Sponsor banner. At this time there is no way to know whether additional sponsors will be Exhibiting Spon sors or Major Sponsors. Therefore, for budgeting purposes you should increase the cost of the banners by the percentage increase in sponsor contributions. (Hint: RoundCA case 5 Cases for Management Decision-Making all calculations to three decimal places.) By checking prices, the Circular Club will be able to obtain insurance providing essentially the same amount of coverage as this year for only $600. For the first rodeo the Club ordered 10,000 tickets. Realizing the corn straints on available seating, the Club is ordering only 5,000 tickets for next year, and therefore its costs are reduced 50%. The sand for the arena for next year will be $300, and miscellaneous fixed costs are to be budgeted at $100 11. A few members in the Circular Club do not want to continue with the annual rodeo. undraiser. Shelley argues that she has spent hundreds of dollars on western ots, hats, and other items of clothing to wear to the rodeo. Are the expenses related However, Shelley is insistent that the Club must continue to conduct the rodeo as an annual f bo ot? to Shelleys purchases of rodeo clothing relevant costs? Why or why r 12. Rather than hire t he local catering company to cater the Contestant Hospitality Tent, members of the Circular Club are considering asking Shadys Bar-B-O to cater the event in exchange for a $600 Major Sponsor spot. In addition, The Fun Shop, a local party will supply business, will be asked to donate a tent to use for the event. The Fun Shop also be given a $600 Major Sponsor spot. Several members of the Club are opposed to this consideration, arguing that the two Major Sponsor spots will take away from the money to be earned throgh other sponsors. Adrian Stein has explained to the mem ers that the Major Sponsor signs for the arena cost only $48 each. In addition, there is more than enough room to display two additional sponsor signs. What would you encourage the Club to do concerning the Contestant Hospitality Tent? Would your an- swer be different if the arena were limited in the number of additional signs that could be displayed? What kind of cost would we consider in this situation that would not be found on a financial statement?
Answer #1

1. Yes, it was necessary for Shelly Jones to mention that she wanted a fundraiser. Because as she took charge as President of the club it was her responsibility to (1) continue to get better each year (2) give back to the society (3) provide the club a presence in the community, for this goals she needed to arrange the fundraiser event.

2. Although the club had made some losses Jonathan said that club had made an investment in the rodeo. Because it can be clearly seen that the club made a presence in the community through the event, in the first year only it had collected significant amount from the sponsors and from ticket sales, it could be a profitable event through little effort from the next year. So it is not a bad investment at all.

3. Yes, Jonathan’s comment is consistent with Shelly’s idea because that is part of what he was meaning by his statement about the club’s investment in the rodeo. Not only he could see it becoming profitable, but also he was looking at it from the perspective that it was benefiting the community and that, in the long run would be something that the club would be known for.

4. Yes, there is a direct relationship of rodeo expenditure and ticket sales because it has incurred expenditure for promotional activities like sponsors sign for arena, western hats for first 500 children. The the more promotional features of the rodeo, the more people will probably want to come. Also there are lots of expenditures too which are directly not related to ticket sales.

5. High Low Method provides an easy way to split fixed and variable components of combined costs using the following formula.

Variable Cost Per Unit= Highest Activity Cost-Lowest Activity Cost/ Highest Activity Units-Lowest Activity Units

Once variable cost per unit is found, you can calculate the fixed cost by subtracting the total variable cost at a specific activity level from the total cost at that activity level.

Fixed Cost Highest Activity Cost - (Variable Cost Per Units x Highest Activity Units)  =  

Or

Fixed Cost Lowest Activity Cost - (Variable Cost Per Units x Lowest Activity Units)  =  

Here, highest activity level is 96 and lowest activity level is 68

And highest activity cost is $1,243 and lowest activity cost is $998

Variable Cost = 1,243-998/96-68 = $8.75

Fixed Cost = 1,243-(8.75*96) = $403

From the calculation we can see that the fixed cost component of the catering costs is $403, and the variable cost component is $8.75 per contestant.

6. Suggested steps to the committee:-

  • Review the expenses incurred in previous year rodeo event to avoid unnecessary expenses

Look for quotation from other life stock contractor to get the cheaper price

  • Ticket must be dated to avoid re-used of ticket by other individual

  • To keep strong gate control

  • Should negotiate with local caterers, hotels, and other local business owners, costs can be cut dramatically

  • Should talk to sponsors who sponsored the rodeo to increase the sponsorship collection

  • Spend little more in promotional activities to increase the ticket sale

7. Variable cost is 4% of gross receipts. So contribution margin is

= 100%-4% = 96%

Given fixed cost = $51,000

BEP in dollars = 51,000/96% = 51,000*100/96 = $53,125

8. (a) Dollars of ticket sale needed in order to earn a target profit of $6000 = Fixed Cost+Target Income/Contribution Margin = $51,000 + $6000 /96% = $59,375

Dollar of ticket sale needed = Required Gross Receipt - Income from Sponsors = 59,375 - 25,600(given in q8) = $33,775

(b) Dollars of ticket sale needed in order to earn a target profit of $12,000 = Fixed Cost+Target Income/Contribution Margin =

$51,000 + $12,000/96% = $65,625

Dollar of ticket sale needed = Required Gross Receipt - Income from Sponsors = 65,625 - 25,600(given in q8) = $40,025

9. Let's calculate the average price of the ticket = $10 + $8 + $6 + $8 / 4 = $8

For target income of $6000 we need $33,775 from ticket sales. So, patrons needed for this = $33,775/$8 = 4,222 patrons (Rounded off)

And the event held for 3 days, so attendance per day required is = 4,222/3 = 1,407 (Rounded off)

Since the arena could hold 2,500 patrons, the facilities at the fairgrounds are adequate to handle crowds.